OpenStudy (anonymous):

If a piece of sea floor has moved 50 km in 5 million years, what is the yearly rate of seafloor motion

OpenStudy (anonymous):

50/5,000,000

OpenStudy (whpalmer4):

the rate is just the total amount moved, divided by the total time taken \[r = \frac{50\text{ km}}{5000000\text{ years}} = \frac{50000 \text{ m}}{5000000\text{ years}} = \text{_______} \frac{\text{ m}}{\text{ year}}\]

OpenStudy (whpalmer4):

If you just write down a number, it will be wrong. A rate is useless without the accompanying units.

OpenStudy (whpalmer4):

Considering the numbers, perhaps cm/year would be an even better unit for the answer.